OpenStudy (bekkah323):

find an equation of the line having the given slope and containing the given point. m=2/3, (1,-2)

OpenStudy (luigi0210):

Use the point slope formula

OpenStudy (luigi0210):

y-y1=m(x-x1)

OpenStudy (bekkah323):

i thought it was y=mx+b

OpenStudy (bekkah323):

or is that the same thing

OpenStudy (luigi0210):

Yes, that's slope-intercept form. You use the point-slope formula to get to slope-intercept form.

OpenStudy (bekkah323):

iv never heard of the point-slope formula, how do i do it?

OpenStudy (luigi0210):

Just plug in values and solve for y. So for this one, it has a slope(m) of 2/3, and the points (1(x1), -2(y1))

OpenStudy (luigi0210):

Now plug all that into here: y-y1=m(x-x1)

OpenStudy (bekkah323):

-2-y1=2/3(1-x1) ?

OpenStudy (luigi0210):

^_^'

OpenStudy (luigi0210):

y-(-2)=2/3(x-1)